3(x+6)=4x I only get as far as 3x+18-4x.I don't know how to finish the problem. Please help
Answer by mananth(16949)

3(x+6)=4x I only get as far as 3x+18-4x.I don't know how to finish the problem. Please help
..
3(x+6)=4x
3x+18=4x
..
bring the variables on one side and the constants on the other side of the equation
add -18
3x+18-18=4x-18
3x=4x-18
add -4x
3x-4x = 4x-4x-18
-x=-18
/-1
x= 18
...
m.ananth@hotmail.ca